# Fairness in Serving Large Language Models

### v1

@@ 5.2 Results on Synthetic Workloads
We start with scenarios where requests arrive deterministically with the same input and output length. In Figure˜3, two clients send requests at different rates, but are both constantly overloaded. In this case, Figure˜3(a) shows VTC can keep the difference between services received by both clients to be small. FCFS cannot maintain fairness, which always serves more for the client who is sending requests at a higher rate. Figure˜3(b) shows the real-time received service rate for two clients in VTC, which confirms that the two received the same level of services at any time interval. This experiment empirically validates Theorem 4.4.
In Figure˜4, three clients send requests at around 2/132/132 / 13, 4/134/134 / 13, and >7/13>7/13> 7 / 13 of the server’s capacity, respectively. In this case, Clients 1 and 2 can be served immediately when their requests arrive (Figure˜4(b)), and Client 3 will consume the remaining capacity (more than 1/3), which is an empirical illustration of the work-conserving property of VTC. The service received for Client 1 and Client 2 have a ratio 1 : 2, which is consistent with their request rates (15 versus 30).
@@
ON/OFF request pattern
In real-world applications, clients usually do not always send requests to the server. They may occasionally be idle (“OFF” phase). We call this the “ON/OFF” pattern. In Figure˜5, Client 2 is always in the “ON” phase, sending requests at a rate of 120 per minute. Client 1 sends 30 requests per minute (less than half of the capacity) during the ON phase and switches to OFF phase periodically. Since Client 1 uses less than half the system capacity when it is in the ON phase, its requests are mostly processed before it switches to the OFF phase (Figure˜5(b)). When it is in the OFF phase, Client 1 thus takes all the system capacity. The total service rate remains the same, which confirms VTC’s flexibility in achieving work-conserving.
On the contrary, in Figure˜6, client 1 sends much more than half the capacity during the ON phase, and makes itself always backlogged. Thus, even when it is in the OFF phase, it is still in the backlog status. In this case, Client 1 and Client 2 should still receive the same level of service rate.
Variable input/output length and poisson process
In this experiment, we simulate scenarios where requests arrive stochastically. Furthermore, they send requests with different input and output lengths. In both Figure˜7 and Figure˜8, Client 1 sends requests with a high rate and Client 2 sends requests with a rate lower but still over its share. Requests arrive according to a Poisson process with the coefficient of variance 1. In Figure˜7, client 1 sends short requests, and client 2 sends long requests. In Figure˜8, Client 1 sends requests with short input and long output, while Client 2 sends requests with long input and short output. Similarly, with the observation before, VTC maintains a bounded difference between the services received by two clients. FCFS cannot preserve fairness according to Figure˜7(b) and Figure˜8(b). This confirms that VTC can work under stochastic workloads with variable lengths.
@@
Isolation
To illustrate the isolation property, we use the setup with a deterministic arrival pattern and the same input length and output length of 256. In Figure˜9, Client 1 sends 30 requests per minute, which is under half of the server’s capacity. Client 2 acts as an "ill-behaved" client. It sends requests at a linearly increasing rate, and gradually over half of the system capacity. We observe that the response time of requests from client 1 is roughly unchanged, empirically validating the property stated in Theorem 4.13.
@@
In reality, clients’ behavior may change over time. To this end, we evaluate the robustness of VTC when the distribution of client requests shifts. In Figure˜10, we construct a 15-minute workload comprising three phases. The first phase is an ON/OFF phase, in which Client 1 sends requests less than its share only during the ON phase and stops during the OFF phase. Client 2 sends requests at a constant rate, which makes the server overloaded. We can observe the pattern for the first phase to be similar to Figure˜5(a), which maintains a constant total service. During the second phase, because the two clients both send requests over their share, a fair server should let them receive the same level of service. Figure˜10(a) demonstrates that VTC yields a desired pattern, similar to that shown in Figure˜3(b). Figure˜10(b) reveals that LCF disproportionately serves Client 1, as it inherits Client 1’s deficit from the first phase. In the last phase, the serving pattern for VTC and LCF are similar, because they simply serve all requests from Client 1 immediately as Client 1 sends requests under its share.
@@ 5.3 Results on Real Workloads
[…]
Figure 11: Response time of FCFS (Left) and VTC (Right) in real traces. Each curve corresponds to one client.
Figure 12: Response time of RPM in real traces. Left to right corresponds to a different rate limit (5, 10, 15, 20 requests per minutes, respectively).
@@
Effect on response time Figure 11 shows the response time of different clients on the real trace. With FCFS scheduling, the response time of all clients increases drastically because some clients send over their share, monopolizing the service and impacting other clients. With VTC, only clients that send requests over its share will have a drastic increase in the response time.
Analysis of request rate per limit approach In Figure˜12, we show the response of RPM approach with different rate limits. In Figure˜13, we show the corresponding throughput comparison with VTC. These plots reveal a core dilemma of the RPM approach - the system has to choose between fairness or throughput, but not both. If the rate limit is low, then the system rejects many requests from clients that send over their share. This opens the capacity for clients with fewer requests. As demonstrated in the uppermost plot, all requests have a similar response time. However, this low rate limit rejects more requests than needed, causing a lower throughput (cluster-wise throughput is 220$\approx$ 220≈ 220 output tokens per second when RPM=5, as opposed to 500$\approx$ 500≈ 500 tokens per second in VTC or FCFS). When the rate limit is set higher, the system throughput is gradually increasing, i.e., increasing from 224 tokens to 450 tokens per second. However, the response time for all requests grows up. When the request rate is set higher and higher, the response time curve converges to the one in FCFS, and there is no fairness guarantee anymore. In other words, the RPM approach can be summarized as follows: it functions as an FCFS (First-Come, First-Served) approach with admission control (rate limiting), rather than as a truly fair scheduler. Its fairness is achieved by rejecting numerous requests from other clients, which compromises the overall system throughput.
@@ 5.4 Ablation Study
In Figure˜14, we evaluate how different memory pool sizes and request lengths will affect scheduling fairness. As shown in Figure˜14(a), with a larger memory pool size, the attainable batch size becomes larger. Therefore, there is greater variation in the absolute difference of accumulated services received by the clients when the memory pool is 65000 than that is 35000, which empirically validates Theorem˜4.4. Figure˜14(b) demonstrates that larger request lengths will also lead to greater variations in the service difference. This is caused by the unknown output length of request generation. At line 19 in Algorithm˜2, the most conservative way of only counting the input tokens leads to over-compensation for the smallest counter, as all the potential output tokens are not counted. A shorter request length has a milder effect of over-compensation. The curves of (5122)(512*2)( 512 ∗ 2 ) and (7682)(768*2)( 768 ∗ 2 ) show the same variance. This is because at length (5122)(512*2)( 512 ∗ 2 ), the upper bound given by VTC has been reached.
@@ 6 Related Works
How to improve the performance of LLM serving systems has recently gained significant attention. Notable techniques cover advanced batching mechanisms [10, 46], memory optimizations [40, 20], GPU kernel optimizations [44, 1, 30, 6], model parallelism [35, 1, 21], parameter sharing [50], and speculative execution [42, 24] were proposed. FastServe [45] explored preemptive scheduling to minimize job completion time (JCT). However, none of these works consider fairness among clients. Our work bridges this gap, and our proposed scheduling methods can be easily integrated with many of these techniques. Our implementation has already integrated with continuous batching (iteration-level scheduling) [46, 16] and PagedAttention [20].
@@ Appendix A Missing Proofs in Proving Fairness of VTC
In Algorithm˜2, $\min_{i\in Q}(c_{i})$ is non-decreasing during the time when $Q\neq\emptyset$.
@@
  • If the condition of line 7 is satisfied, at lines 8-14, a new client will be added to $Q$. If lines 10-11 are reached, the $\min_{i\in Q}(c_{i})$ is equals to its value at the last time when $Q\neq\emptyset$. If line 13 is reached, since $c_{u}=\max\{c_{u},\min_{i\in Q}c_{i}\}$, the $\min_{i\in Q}(c_{i})$ will not change.
  • At line 23 and line 29, the $c_{i}$’s can only increase, so that $\min_{i\in Q}(c_{i})$ is non-decreasing.
  • At line 25, if a client has cleared all its requests from $Q$, that the client is removed from $Q$, $\min_{i\in Q}(c_{i})$ cannot decrease.
@@
We prove the lemma by induction. During the induction, for each line of change of $c_{i}$ in Algorithm˜2, we use $c_{i}^{\prime}$ to denote the new value and $c_{i}$ to denote the original value. Similarly, we use $Q^{\prime}$ to donate the new value and $Q$ to denote the original value. We also use $c_{i}^{(t)}$ to denote the value of $c_{i}$ at time $t$, and $Q^{(t)}$ to denote the value of $Q$ at time $t$.
@@
    If a client $u\notin Q$ receive a new request and thus $Q^{\prime}=Q\cup\{u\}$, line 8-9 will be reached, and thus $c^{\prime}_{u}=\max\{c_{u},\min_{i\in Q}c_{i}\}\geq\min_{i\in Q}c_{i}$. Then we have,
@@
    Let $t$ be the last time that $u$ was in $Q$ before the change, and thus $c_{u}^{(t)}=c_{u}$. From Equation˜2, there is
@@
    From Lemma˜A.1, there is $\min_{i\in Q^{(t)}}c_{i}^{(t)}\leq\min_{i\in Q}c_{i}$, so we have
@@
    Combine with Equation˜2 and Equation˜4, there is
@@
    Therefore, Equation˜2 holds after the change.
@@
    If a client $u$ is left from $Q$ at line 21, the difference $\max_{i\in Q}c_{i}-\min_{i\in Q}c_{i}$ will not increase. Because $\max(C^{\prime})-\min(C^{\prime})\leq\max(C)-\min(C),\forall C\supseteq C^{\prime},C^{\prime}\neq\emptyset$. Therefore, Equation (2) still holds.
@@
    At line 19, since $c_{k}=\min_{i\in Q}c_{i}$, there is
@@
    In Equation˜5 we have derived that:
@@
    Thus Equation˜6 gives:
@@
    Finally, combining the inequality from Equation˜5 that
@@
    Therefore, Equation˜2 holds.
@@
    At line 25, let $k=\operatorname*{arg\,max}_{i\in Q}c_{i}^{\prime}$, so that $c_{k}^{\prime}=\max_{i\in Q}c_{i}^{\prime}$. Let $r$ be the last one among requests from $k$ that have been scheduled. Let $t$ be the time when $r$ was selected at line 16. Since $r$ is the last one been scheduled from $k$, there is
@@
    Because request $r$ from client $k$ has been scheduled at time $t$, from line 16, there is $c_{k}^{(t)}=\min_{i\in Q^{(t)}}c_{i}^{(t)}$. From Lemma A.1, we have $\min_{i\in Q^{(t)}}c_{i}^{(t)}\leq\min_{i\in Q}c^{\prime}_{i}$. Combine with Equation˜8, we have
@@
    Therefore, Equation˜2 holds.
@@
Consider at time 0 the client $f$ sends a list of requests which cannot fit in the memory at once. Because of work-conserving, client $f$ will fill the whole running batch. In this case, client $f$ is backlogged, and any new query is not processed until the existing queries finish processing. Assume that all existing queries finish at time $T$, and that at time $\epsilon$ with $\epsilon$ close to 0, a second client $g$ sends another batch of requests. Now during the time interval $[\epsilon,T]$, both clients $f,g$ are backlogged since there exist queries from both clients in the queue. At time $T$, client $f$ received service from the first batch of processing, which can be up to $w_{q}\cdot M$ if the memory is luckily fully utilized. Thus we have
@@
Let the counter for $f$ be $c_{f}$ after line 9 for $r_{f}$. Before $D_{r_{f}}$, since $r_{f}$ is always in the queue, the counter for f will not be lifted. Since there is no running batch of $f$ in the server, line 17 will select $r_{f}$ to be the next one for $f$. Lemma 4.3 shows that for any other client g,
@@
Since $c_{i}$’s in Algorithm˜2 are non-decreasing,
@@
Combine with Equation˜9, there is:
@@
From the pigeonhole principle, there is at least one client $g$ who has received services $W_{g}(t,t_{2})\geq\frac{T(t,t_{2})}{n(t,t_{2})}$. If $f=g$, the claim holds. If not, from Theorem˜4.9, we have

### v2

@@ 5.2 Results on Synthetic Workloads
We start with scenarios where requests arrive deterministically with the same input and output length. In Figure 3, two clients send requests at different rates, but are both constantly overloaded. In this case, Figure 3(a) shows VTC can keep the difference between services received by both clients to be small. FCFS cannot maintain fairness, which always serves more for the client who is sending requests at a higher rate. Figure 3(b) shows the real-time received service rate for two clients in VTC, which confirms that the two received the same level of services at any time interval. This experiment empirically validates Theorem 4.4.
In Figure 4, three clients send requests at around $2/13$, $4/13$, and $>7/13$ of the server’s capacity, respectively. In this case, Clients 1 and 2 can be served immediately when their requests arrive (Figure 4(b)), and Client 3 will consume the remaining capacity (more than 1/3), which is an empirical illustration of the work-conserving property of VTC. The service received for Client 1 and Client 2 have a ratio 1 : 2, which is consistent with their request rates (15 versus 30).
ON/OFF request pattern
In real-world applications, clients usually do not always send requests to the server. They may occasionally be idle (“OFF” phase). We call this the “ON/OFF” pattern. In Figure 5, Client 2 is always in the “ON” phase, sending requests at a rate of 120 per minute. Client 1 sends 30 requests per minute (less than half of the capacity) during the ON phase and switches to OFF phase periodically. Since Client 1 uses less than half the system capacity when it is in the ON phase, its requests are mostly processed before it switches to the OFF phase (Figure 5(b)). When it is in the OFF phase, Client 1 thus takes all the system capacity. The total service rate remains the same, which confirms VTC’s flexibility in achieving work-conserving.
On the contrary, in Figure 6, client 1 sends much more than half the capacity during the ON phase, and makes itself always backlogged. Thus, even when it is in the OFF phase, it is still in the backlog status. In this case, Client 1 and Client 2 should still receive the same level of service rate.
@@
Variable input/output length and poisson process
In this experiment, we simulate scenarios where requests arrive stochastically. Furthermore, they send requests with different input and output lengths. In both Figure 7 and Figure 8, Client 1 sends requests with a high rate and Client 2 sends requests with a rate lower but still over its share. Requests arrive according to a Poisson process with the coefficient of variance 1. In Figure 7, client 1 sends short requests, and client 2 sends long requests. In Figure 8, Client 1 sends requests with short input and long output, while Client 2 sends requests with long input and short output. Similarly, with the observation before, VTC maintains a bounded difference between the services received by two clients. FCFS cannot preserve fairness according to Figure 7(b) and Figure 8(b). This confirms that VTC can work under stochastic workloads with variable lengths.
@@
Isolation
To illustrate the isolation property, we use the setup with a deterministic arrival pattern and the same input length and output length of 256. In Figure 9, Client 1 sends 30 requests per minute, which is under half of the server’s capacity. Client 2 acts as an "ill-behaved" client. It sends requests at a linearly increasing rate, and gradually over half of the system capacity. We observe that the response time of requests from client 1 is roughly unchanged, empirically validating the property stated in Theorem 4.13.
@@
In reality, clients’ behavior may change over time. To this end, we evaluate the robustness of VTC when the distribution of client requests shifts. In Figure 10, we construct a 15-minute workload comprising three phases. The first phase is an ON/OFF phase, in which Client 1 sends requests less than its share only during the ON phase and stops during the OFF phase. Client 2 sends requests at a constant rate, which makes the server overloaded. We can observe the pattern for the first phase to be similar to Figure 5(a), which maintains a constant total service. During the second phase, because the two clients both send requests over their share, a fair server should let them receive the same level of service. Figure 10(a) demonstrates that VTC yields a desired pattern, similar to that shown in Figure 3(b). Figure 10(b) reveals that LCF disproportionately serves Client 1, as it inherits Client 1’s deficit from the first phase. In the last phase, the serving pattern for VTC and LCF are similar, because they simply serve all requests from Client 1 immediately as Client 1 sends requests under its share.
@@ 5.3 Results on Real Workloads
We construct real workload traces from the traces of LMSYS Chatbot Arena [53, 52], following a similar process in [43]. The trace is from a server that serves multiple LLMs. To adapt it to our setting, we treat each LLM as a client. In total, there are 27 clients. To sample from this log, we define $D$, the duration, and $R$, the request rate. We then sample $R*D$ requests from the trace, and re-scale the real-time stamps to $[0,D]$. We use a duration of 10 minutes to be consistent with previous experiments, and a request rate of 210 requests per minute for the whole system. With the adapted workload, we run Llama-2-7b on A10G (24GB). In summary, the prompts from the 27 clients are collected from the real world interactions, which will be sent to the server for inference on Llama-2-7b. The timestamps are re-scaled from the real-world trace.
For better visualization of the evaluation results, we select two clients that send the most requests and two clients that send a medium number of requests. We sort 27 clients according to the number of requests they send, and depict the statistics of the $13^{th},14^{th}$ and $26^{th},27^{th}$ clients. We do not choose clients that send the least requests because they typically only send requests in a small interval.
Figure 11: Request rate distribution during the sampled 10 minutes duration with re-scale. The figure on the left denotes the real-time request rate for the 27 clients. A few clients have sent many more requests than others, reflecting the original trace of a few most popular models. The figure on the right depicts the total request rate from all 27 clients.
Figure 12: Response time of 4 selected clients when using FCFS (Left) and VTC (Right) in real traces. Each curve corresponds to one client. There are some curves that show disconnected because, during some periods, a client may have no requests served. Requests distribution see Figure 11.
Figure 13: Response time of 4 selected clients (all 27 clients when rpm=5) when using RPM in real traces. Left-upper to right-bottom corresponds to a different rate limit (5, 15, 20, 30 requests per minutes, respectively). There are some curves that show disconnected because, during some periods, a client may have no requests served.
@@
Request distribution The request rate distribution is visualized in Figure 11. The request rate of individual clients and the total request rate are all highly dynamic. The input and output length distribution is depicted in Figure 20 in the Appendix. The average input length is 136, and the average output length is 256. The input and output lengths have the range of $[2,1021]$ and $[2,977]$, respectively.
Effect on response time Figure 12 shows the response time of 4 selected clients on the real trace. With FCFS scheduling, the response time of all clients increases drastically because some clients send over their share, monopolizing the service and impacting other clients. With VTC, only clients that send requests over its share will have a drastic increase in the response time.
Analysis of request rate per limit approach In Figure 13, we show the response of RPM approach with different rate limits. In Figure 14, we show the corresponding throughput comparison with VTC. These plots reveal a core dilemma of the RPM approach - the system has to choose between fairness or throughput, but not both. If the rate limit is low, then the system rejects many requests from clients that send over their share. This opens the capacity for clients with fewer requests. As demonstrated in the uppermost plot in Figure 13, all requests have a similar response time. However, this low rate limit rejects more requests than needed, causing a lower throughput (cluster-wise throughput is $\approx 340$ output tokens per second when RPM=5, as opposed to $\approx 779$ tokens per second in VTC or FCFS). When the rate limit is set higher, the system throughput is gradually increasing, i.e., increasing from $340$ tokens to $747$ tokens per second. However, the response time for all requests grows up. When the request rate is set higher and higher, the response time curve converges to the one in FCFS, and there is no fairness guarantee anymore. In other words, the RPM approach can be summarized as follows: it functions as an FCFS (First-Come, First-Served) approach with admission control (rate limiting), rather than as a truly fair scheduler. Its fairness is achieved by rejecting numerous requests from other clients, which compromises the overall system throughput.
Quantitative Measurement We measured the maximum and average service difference described in Section 5.1 during the time window (10 minutes) in which we ran the experiments. Table 2 is a summary for all baselines using this quantitative measurement for real workload trace.
Scheduler Max Diff Avg Diff Diff Var Throu Isolation
FCFS 759.97 433.53 32112.00 777 No
LCF 750.49 323.82 29088.90 778 Some [^9] LCF achieves isolation if the workload does not change. However, the isolation can be broken by newly joined clients whose virtual counter is lagging behind.
VTC 368.40 251.66 6549.16 779 Yes
VTC(predict) 365.47 240.33 5321.62 773 Yes
VTC(oracle) 329.46 227.51 4475.76 781 Yes
RPM(5) 143.86 83.58 1020.46 340 Some
RPM(20) 446.76 195.71 7449.79 694 Some
RPM(30) 693.66 309.45 24221.31 747 Some
Table 2: The service difference is counted by summing the service difference between each client and the client who received the maximum services. Throughput is the total number of tokens (including input and output tokens) processed divided by the total execution time.
@@ 5.4 Ablation Study
In Figure 15, we evaluate how different memory pool sizes and request lengths will affect scheduling fairness. As shown in Figure 15(a), with a larger memory pool size, the attainable batch size becomes larger. Therefore, there is greater variation in the absolute difference of accumulated services received by the clients when the memory pool is 65000 than that is 35000, which empirically validates Theorem 4.4. Figure 15(b) demonstrates that larger request lengths will also lead to greater variations in the service difference. This is caused by the unknown output length of request generation. At line 24 in Algorithm 2, the most conservative way of only counting the input tokens leads to over-compensation for the smallest counter, as all the potential output tokens are not counted. A shorter request length has a milder effect of over-compensation. The curves of $(512*2)$ and $(768*2)$ show the same variance. This is because at length $(512*2)$, the upper bound given by VTC has been reached.
@@ 6 Related Works
How to improve the performance of LLM serving systems has recently gained significant attention. Notable techniques cover advanced batching mechanisms [13, 50], memory optimizations [44, 24], GPU kernel optimizations [48, 2, 34, 9], model parallelism [39, 2, 25], parameter sharing [55], and speculative execution [46, 28] were proposed. FastServe [49] explored preemptive scheduling to minimize job completion time (JCT). However, none of these works consider fairness among clients. Our work bridges this gap, and our proposed scheduling methods can be easily integrated with many of these techniques. Our implementation used for this paper is built atop continuous batching (iteration-level scheduling) [50] [^10] We presented VTC on continuous batching with separated prefill and decode steps in the main context. A general integration is discussed in Section C.1. and PagedAttention [24].
@@ Appendix A Missing Proofs in Proving Fairness of VTC
In Algorithm 2, $\min_{i\in Q}(c_{i})$ is non-decreasing during the time when $Q\neq\emptyset$.
@@
  • If the condition of line 7 is satisfied, at lines 8-14, a new client will be added to $Q$. If lines 9-10 are reached, the $\min_{i\in Q}(c_{i})$ is equals to its value at the last time when $Q\neq\emptyset$. If lines 12-13 are reached, since $c_{u}=\max\{c_{u},\min_{i\in Q}c_{i}\}$, the $\min_{i\in Q}(c_{i})$ will not change.
  • At line 24 and line 30, the $c_{i}$’s can only increase, so that $\min_{i\in Q}(c_{i})$ is non-decreasing.
  • At line 26, if a client has cleared all its requests from $Q$, that the client is removed from $Q$, $\min_{i\in Q}(c_{i})$ cannot decrease.
@@
We prove the lemma by induction. During the induction, for each line of change of $c_{i}$ in Algorithm 2, we use $c_{i}^{\prime}$ to denote the new value and $c_{i}$ to denote the original value. Similarly, we use $Q^{\prime}$ to donate the new value and $Q$ to denote the original value. We also use $c_{i}^{(t)}$ to denote the value of $c_{i}$ at time $t$, and $Q^{(t)}$ to denote the value of $Q$ at time $t$.
@@
    If a client $u\notin Q$ receive a new request and thus $Q^{\prime}=Q\cup\{u\}$, line 12-13 will be reached, and thus $c^{\prime}_{u}=\max\{c_{u},\min_{i\in Q}c_{i}\}\geq\min_{i\in Q}c_{i}$. Then we have,
@@
    Let $t$ be the last time that $u$ was in $Q$ before the change, and thus $c_{u}^{(t)}=c_{u}$. From Equation 2, there is
@@
    From Lemma A.1, there is $\min_{i\in Q^{(t)}}c_{i}^{(t)}\leq\min_{i\in Q}c_{i}$, so we have
@@
    Combine with Equation 2 and Equation 4, there is
@@
    Therefore, Equation 2 holds after the change.
@@
    If a client $u$ is left from $Q$ at line 26, the difference $\max_{i\in Q}c_{i}-\min_{i\in Q}c_{i}$ will not increase. Because $\max(C^{\prime})-\min(C^{\prime})\leq\max(C)-\min(C),\forall C\supseteq C^{\prime},C^{\prime}\neq\emptyset$. Therefore, Equation (2) still holds.
@@
    At line 24, since $c_{k}=\min_{i\in Q}c_{i}$, there is
@@
    In Equation 5 we have derived that:
@@
    Thus Equation 6 gives:
@@
    Finally, combining the inequality from Equation 5 that
@@
    Therefore, Equation 2 holds.
@@
    At line 30, let $k=\operatorname*{arg\,max}_{i\in Q}c_{i}^{\prime}$, so that $c_{k}^{\prime}=\max_{i\in Q}c_{i}^{\prime}$. Let $r$ be the last one among requests from $k$ that have been scheduled. Let $t$ be the time when $r$ was selected at line 21. Since $r$ is the last one been scheduled from $k$, there is
@@
    Because request $r$ from client $k$ has been scheduled at time $t$, from line 20, there is $c_{k}^{(t)}=\min_{i\in Q^{(t)}}c_{i}^{(t)}$. From Lemma A.1, we have $\min_{i\in Q^{(t)}}c_{i}^{(t)}\leq\min_{i\in Q}c^{\prime}_{i}$. Combine with Equation 8, we have
@@
    Therefore, Equation 2 holds.
@@
Consider at time 00 the client $f$ sends a list of requests which cannot fit in the memory at once. Because of work-conserving, client $f$ will fill the whole running batch. In this case, client $f$ is backlogged, and any new query is not processed until the existing queries finish processing. Assume that all existing queries finish at time $T$, and that at time $\epsilon$ with $\epsilon$ close to 0, a second client $g$ sends another batch of requests. Now during the time interval $[\epsilon,T]$, both clients $f,g$ are backlogged since there exist queries from both clients in the queue. At time $T$, client $f$ received service from the first batch of processing, which can be up to $w_{q}\cdot M$ if the memory is luckily fully utilized. Thus we have
@@
Let the counter for $f$ be $c_{f}$ after line 13 for $r_{f}$. Before $D_{r_{f}}$, since $r_{f}$ is always in the queue, the counter for f will not be lifted. Since there is no running batch of $f$ in the server, line 21 will select $r_{f}$ to be the next one for $f$. Lemma 4.3 shows that for any other client g,
@@
Since $c_{i}$’s in Algorithm 2 are non-decreasing,
@@
Combine with Equation 9, there is:
@@
From the pigeonhole principle, there is at least one client $g$ who has received services $W_{g}(t,t_{2})\geq\frac{T(t,t_{2})}{n(t,t_{2})}$. If $f=g$, the claim holds. If not, from Theorem 4.9, we have